OpenStudy (anonymous):

Find the solution set of the equation. 2e^(3x-1) = 7 I understand the problem but the answer is : { 1/3 + 1/3ln(7/2) }

Directrix (directrix):

Take natural logs of both sides of the equation to begin: 2e^(3x-1) = 7 after dividing by 2. e^(3x-1) = 7/2 ln(e^(3x-1) = ln(7/2) What comes next?

OpenStudy (anonymous):

Got it, thanks!!

Directrix (directrix):

Did you want to get the exact answer or an approximation? Just checking.

OpenStudy (anonymous):

I just needed the solution set so { 1/3 + 1/3ln7/2 } works